OpenStudy (dls):

A body of mass 2 kg, moving on a horizontal surface with an initial velocity of 4 m/s, comes to rest after 2 seconds. If one wants to keep this body moving on the same surface with a velocity of 4 m/s, the force required is

OpenStudy (mayankdevnani):

is answer is 4N

OpenStudy (dls):

how

OpenStudy (dls):

yes

OpenStudy (mayankdevnani):

first, find acceleration of body is what is the formula of acceleration of a body???

OpenStudy (dls):

a=v-u/t

OpenStudy (dls):

arent v & u both4?

OpenStudy (mayankdevnani):

no?

OpenStudy (mayankdevnani):

\[a=\frac{v-u}{t}\] where u=4 m/s and v=0 m/s

OpenStudy (mayankdevnani):

find accleration?

OpenStudy (dls):

why? they are saying initial and final velocities should be same..i.e 4m/s it shouldnt change

OpenStudy (mayankdevnani):

no, your question said that initial velocity is 4 m/s( written in your question) and final velocity is 0 m/s(comes to rest) . ok @DLS

OpenStudy (mayankdevnani):

Given u=4 m/s v=0 m/s t=2 s a=v-u/t a=0-4/2 so find acceleration? can you @DLS

OpenStudy (mayankdevnani):

@DLS

OpenStudy (dls):

oh dw I gt it!!

OpenStudy (mayankdevnani):

ok